How can I fin Radius and Height of cylinder when I was provided with circumference and volume in liters.?
Please explain with diagramatic representation. Thanks in advanced

Answer by lwsshak3(11628)

How can I fin Radius and Height of cylinder when I was provided with circumference and volume in liters.?
**
V=volume (given)
A=area of cross-section
h=height of cylinder (to find)
c=circumference(given)
r=radius (to find)
..
V=Ah
h=V/A
A=πr^2
c=2πr
r=c/2π
..
A=πr^2=π(c/2π)^2=π(c^2/4π^2=c^2/4π
h=V/A=V/c^2/4π=4πV/c^2
r=c/2π
ans:
height=4πV/c^2 cm
radius=c/2π cm